you and your friend share a paper route. Youcan deliver 4 papers per minute and your friend can deliver 3 papers per minute. 75 papers have to be delivered each day on the route. Let n be the number of papers you deliver.
a. Use the verbal model to write an equation that you can use to findout how long it will take the both of you togeter to deliver the papers.
total amount of time = your rate*number of papers you deliver + friends
rate*number of papers friend delivers




b.How long will it take the both of you to deliver the papers if you deliver 38 papers? 50 papers?

Answer by checkley77(12569)

Every minute you & your friend deliver 3+4=7 papers.
Therefoe you divide the number of papers to bw delivered by 7 to find out how long it will take.
75/(4+3)=75/7=10.7 minutes to deliver the papers together.
------------------------------------------------------------
Sorry I missed the wording on the next questions.
(75-38)/3=37/3=12.333. minutes for your friend to deliver his 37 papers.
(75-50)/3=25/3=8.333 min. for your friend to deliver his 25 papers.